The area of a trapezoid is A=(a+b)h over 2. If A=300, h=5, and b=30, what is the value of a?
Answer by rfer(16322)

A=(a+b)h
a=A/h-b
a=300/5-30
a=60-30
a=30
Sorry for the missunderstanding, I did not see the over 2 before.
A=((a+b)h)/2
2A=ah+bh
ah=2A-bh
a=(2A-bh)/h
a=(2(300)-(30)(5))/5
a=(600-150)/5
a=450/5=90
I hope that looks better to you.
Again, sorry.
Bob